What is the value of the system determinant for the following system of equations?
6x + y = 4
5x + y - 7 = 0

Found 2 solutions by drj, Alan3354:
Answer by drj(1380)

What is the value of the system determinant for the following system of equations?
6x + y = 4
5x + y - 7 = 0 or 5x+y=7.

The determinant for this system of equations is 6*1-5*1=1.

I hope the above steps were helpful.

For FREE Step-By-Step videos in Introduction to Algebra, please visit
http://www.FreedomUniversity.TV/courses/IntroAlgebra and for Trigonometry visit
http://www.FreedomUniversity.TV/courses/Trigonometry.

Good luck in your studies!

Respectfully,
Dr J

Answer by Alan3354(21608)

What is the value of the system determinant for the following system of equations?
6x + y = 4
5x + y - 7 = 0
---------
It's 6*1 - 5*1 = 1